# Learning Decentralized Multi-Biped Control for Payload Transport

### v1

@@ 1 Introduction
Multi-biped transport presents significant challenges. Legged robots are generally unstable and highly dynamic, making them much harder to control compared to wheeled robots. Maintaining stability while transporting payloads and coordinating multiple legged-robots adds further complexity to the control problem [6]. Prior works on multi-legged transport systems have primarily focused on model-based approaches that rely on accurate modeling of robot interaction and explicit inter-robot communication, which limit their reconfigurability and practicality [7, 8]. While learning-based approaches have emerged as promising alternatives [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] for various control problems, existing methods on multi-robot system often assume centralized control with strong assumptions on system observability, leading to limited scalability, robustness, and vulnerability to point of failure [19].
@@ 4.3 Reward Design
Our reward function is designed to produce stable locomotion behavior while ensuring the robots collaboratively follow the commanded motion of the carrier control point. For each robot $i$ in N-robot configuration, where $i\in\{1,\ldots,N\}$, the reward $r_{i}$ is composed of two components: the local robot reward $r^{L}_{i}$ and global reward $r^{G}$, where $r_{i}=r^{L}_{i}+r^{G}$ at each timestep throughout the episode. The local reward $r^{L}_{i}$ is designed to produce stable locomotion behavior of each robot, including maintaining consistent base height, penalizing jerky motion, regularizing proper foot alignment and the stepping frequency, with single-contact reward as proposed by van Marum et al. [30]. Furthermore, the local reward minimizes torque of the motors to penalize excessive energy expenditure as well as minimizes the force exerted at the attachment point at the carrier during standing, ensuring that robots do not pull or push against each other but hold the carrier steady. The global reward $r^{G}$ focuses on the performance of the combined action of all robots in following the carrier commands. It aims to minimize the deviation between the commanded and actual x-velocity, y-velocity, angular yaw velocity, and height of the carrier control point. We detail the reward components in Appendix B.
@@ 5.3 Sim-to-real Transfer
We evaluated our policy on real hardware consisting of two and three bipedal robots, Cassie, attached at the endpoints of an I-shaped and T-shaped carrier made from aluminum extrusion with a weight holder. We also added a payload on the weight holder and used it as a carrier control point [Figure 1 (D)]. We deployed the decMBC policy on each of the robots, running independently, providing their relative positions with respect to the carrier control point. The carrier commands were provided using a joystick which would broadcast the same carrier command values to all the robots using UDP protocol. We commanded x-velocity, y-velocity, angular yaw velocity, and height to maneuver the payload from one point to another. We observed that it was successfully able to maneuver different payloads according to various user commands for an extended period of operation.
@@ 6 Limitations
Despite the promising results, there are clear limitations to address in future work. First, the decMBC is not explored on diverse terrain types beyond flat surfaces, eliminating the understanding of its performance in more challenging environments. Second, real-world experiments are demonstrated to up to three Cassie robots, despite the simulation results showing the scalability of up to 10 robots. Third, the controller relies solely on proprioceptive sensing and relative robot positions without incorporating additional sensing modalities or addressing autonomous navigation and high-level decision-making challenges. Integrating the decMBC with perception, planning, and control modules could enhance its capabilities in complex environments. Finally, although our policy is trained with a wide range of relative positions of each robot with respect to the carrier point, which should cover most real-world scenarios, our policy does not guarantee it may work for positions beyond the range we have set during training.

### v2

@@ 1 Introduction
Multi-biped transport presents significant challenges. Legged robots are generally unstable and highly dynamic, making them much harder to control compared to wheeled robots. Maintaining stability while transporting payloads and coordinating multiple legged-robots adds further complexity to the control problem [6]. However, the use of multiple bipeds offers distinct advantages over alternatives such as large quadrupeds, providing greater flexibility in payload transport. This approach allows the number and arrangement of robots to be tailored based on the weight and size of the payload. Such versatility opens up potential applications in various areas including construction, moving systems, warehouses, agriculture, and delivery systems.
Prior works on multi-legged transport systems have primarily focused on model-based approaches that rely on accurate modeling of robot interaction and explicit inter-robot communication, which limit their reconfigurability and practicality [7, 8]. While learning-based approaches have emerged as promising alternatives [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] for various control problems, existing methods on multi-robot system often assume centralized control with strong assumptions on system observability, leading to limited scalability, robustness, and vulnerability to point of failure [19].
@@ 4.3 Reward Design
Our reward function is designed to produce stable locomotion behavior while ensuring the robots collaboratively follow the commanded motion of the carrier control point. For each robot $i$ in N-robot configuration, where $i\in\{1,\ldots,N\}$, the reward $r_{i}$ is composed of two components: the local robot reward $r^{L}_{i}$ and global reward $r^{G}$, where $r_{i}=r^{L}_{i}+r^{G}$ at each timestep throughout the episode. The local reward $r^{L}_{i}$ is designed to produce stable locomotion behavior of each robot, including maintaining consistent base height, penalizing jerky motion, regularizing proper foot alignment and the stepping frequency, with single-contact reward as proposed by van Marum et al. [30]. Furthermore, the local reward minimizes torque of the motors to penalize excessive energy expenditure and minimizes the force exerted at the attachment point at the carrier during standing, ensuring that robots do not pull or push against each other but hold the carrier steady. The global reward $r^{G}$ focuses on the performance of the combined action of all robots in following carrier commands. It aims to minimize the deviation between the commanded and actual x-velocity, y-velocity, angular yaw velocity, and height of the carrier control point. We detail the reward components in Appendix B.
@@ 5.2 Comparison with Centralized and Specialized policies
While our decentralized approach shows comparable or better performance in many scenarios, centralized policies theoretically have higher potential capacity due to their access to global information. However, our attempts at training centralized controllers have not yielded substantial improvements, possibly because our decentralized architecture is sufficiently rich for near-optimal performance under rigid attachments, or because the higher dimensionality of the centralized approach makes it more challenging to train effectively. For instance, training a centralized controller for a 3-robot configuration took about 1000 hours to reach peak reward, compared to 245 hours for our decentralized policy. This suggests that while centralized policies may have higher theoretical capacity, our decentralized approach offers a favorable trade-off between performance and training efficiency.
@@ 5.3 Sim-to-real Transfer
We evaluated our policy on real hardware consisting of two and three bipedal robots, Cassie, attached at the endpoints of an I-shaped and T-shaped carrier made from aluminum extrusion with a weight holder. We also added a payload on the weight holder and used it as a carrier control point [Figure 1 (D)]. We deployed the decMBC policy on each of the robots, running independently, providing their relative positions with respect to the carrier control point. We use a joystick to broadcast the same carrier command values to all the robots using the UDP protocol. We commanded x-velocity, y-velocity, angular yaw velocity, and height to maneuver the payload from one point to another. We observed that it was successfully able to maneuver different payloads according to various user commands for an extended period of operation.
@@ 6 Limitations
Despite the promising results, there are clear limitations to address in future work. First, the decMBC is not explored on diverse terrain types beyond flat surfaces, eliminating the understanding of its performance in more challenging environments. Second, real-world experiments are demonstrated to up to three Cassie robots, despite the simulation results showing the scalability of up to 10 robots. Third, the controller relies solely on proprioceptive sensing and relative robot positions without incorporating additional sensing modalities like camera perception which could enhance its capabilities in complex environments. Fourth, there are several practical challenges associated with multi-biped systems that require further consideration, including higher costs, increased points of failure, perception difficulties due to robot proximity, and the complexity of deployment and maintenance. These factors highlight important trade-offs between system flexibility and operational practicality. Finally, although our policy is trained with a wide range of relative positions of each robot with respect to the carrier point, which should cover most real-world scenarios, our policy does not guarantee it may work for positions beyond the range we have set during training.
Acknowledgments
This material is based on work supported by the National Science Foundation under Grants 2321851 and IIS-1724360 and USDA 2021-67021-35344. We thank all the reviewers for their helpful feedback, which improved the quality of this paper. In addition to the authors, we also appreciate the support from our colleagues, Pranay Dugar and Bart van Marum, for their assistance with robot operation and repair.